jogger runs 202 meters westward, then 176 meters eastward, then 94 meters westward, and finally 312 meters eastward.
For this motion, what is the distance moved?
What is the magnitude and direction of the displacement?
1
Expert's answer
2020-04-01T10:00:06-0400

Distance is


"d=202+176+94+312=784\ m"

Displacement:


"s=-202+176-94+312=192\ m"

Direction: eastward.